An iron has a power of 1000 W. How much energy will it electrically transfer each second?
1
Expert's answer
2020-03-19T11:39:50-0400

Power is the rate of doing work or of transferring heat per unit of time, or, in other words, power is the rate of transferred energy per unit of time


"P=\frac{E}{t}"

"P" - power

"E" - transferred energy

"t" - time it took to tranfer energy "E"


In our case, "P=1000 W, t=1s \rArr E=Pt=1000*1=1000 J"